the length of a rectangle is 9 cm longer than its width. If the total perimeter is 250 cm determine the dimensions of the rectangle
Answer by macston(5194)

width=W, length=L=W+9
Perimeter P=2(L+W)=250 cm substitute for L
P=2(W+9+W)
250 cm=2(2W+9) divide both sides by 2
125 cm=2W+9 subtract 9 from each side
116 cm=2W divide each side by 2
58 cm=W, ANSWER 1 Width=58 cm
L=W+9=58 cm + 9cm= 67 cm ANSWER 2 Length = 67 cm
CHECK
perimeter=2(L+W)
250 cm=2(58cm+67 cm)
250 cm=2(125 cm)
250 cm=250 cm